OpenStudy (anonymous):

The area of a regular octagon varies directly with the square of the length of one of its sides. If an octagon with area 28 square feet has a side length of 2 feet, what is the constant of variation?

OpenStudy (tad1):

A = k (L^2) 28 = k(4) k = 7